Walter is 3 times as old as Marti. Ten years from now, walter will be twice as old as martin will be then. How old is each now?
Answer by ewatrrr(24785)


Hi,
Walter is 3 times as old as Marti
Let x and 3x represent Marti's and Walter's ages NOW respectivey
Write as we Read*** ( ten years from now)
3x + 10 = 2(x+10)
solving for x
3x + 10 = 2x + 20
x = 10, Marti's age NOW. Walter is NOW 30. (3*10)
CHECKING Our Answer****
40 = 2(10+10) = 40